OpenStudy (sugarmochi):

Solve the equation x^2 /2 = x+4 graphically and find the coordinates of their points of intersection algebraically.

OpenStudy (amorfide):

when you have simultaneous equations such as \[y=\frac{ x^{2} }{ 2 }\] \[y=x+4\] since they are both equal to y, we can write it like this \[\frac{ x^{2} }{ 2 }=x+4\] now that we know this we can split what you have, into two equations in the form of y=...

OpenStudy (amorfide):

so you would draw the graph for the two equations above, on the same graph

OpenStudy (amorfide):

your next step, is to just check your grpah that you drew, and get the points of intersection by reading directly from the graph

OpenStudy (amorfide):

to solve algebraically you would use what you are given x^2 /2 = x+4 and solve for x you will end up with 2 values of x, due to having a quadratic then you would split it up again y=x+4 y=x^2/2 then substitute your values of x, into one of these equations, and then you will have two y values, and your points of intersections will be written as \[(x_{1},y_{1}),(x_{2},y_{2})\]

OpenStudy (amorfide):

obviously they will be the values that you worked out

OpenStudy (amorfide):

@SugarMochi are you still there?

Directrix (directrix):

@SugarMochi Is there supposed to be a y-variable? > the equation x^2 /2 = x+4

OpenStudy (zzr0ck3r):

They want the user to solve it graphically. You can think of "them" (left and right of equals) as functions in the form y= , and then solve as stated above.